OpenStudy (anonymous):

help: 4/5x-3.3/4x=2.1/2

OpenStudy (lgbasallote):

\[\frac 45 x - \frac{3.3}{4}x = \frac{2.1}{2}\] right?

OpenStudy (anonymous):